### v1

# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

Abstract.
Recent advancements in image synthesis, particularly with the advent of GAN and Diffusion models, have amplified public concerns regarding the dissemination of disinformation. To address such concerns, numerous AI-generated Image (AIGI) Detectors have been proposed and achieved promising performance in identifying fake images. However, there still lacks a systematic understanding of the adversarial robustness of AIGI detectors. In this paper, we examine the vulnerability of state-of-the-art AIGI detectors against adversarial attack under white-box and black-box settings, which has been rarely investigated so far. For the task of AIGI detection, we propose a new attack containing two main parts. First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution. Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models, e.g. transferring adversarial examples across CNNs and ViTs. This is achieved by introducing a novel post-train Bayesian strategy that turns a single surrogate into a Bayesian one, capable of simulating diverse victim models using one pre-trained surrogate, without the need for re-training. We name our method as frequency-based post-train Bayesian attack, or FPBA. Through FPBA, we show that adversarial attack is truly a real threat to AIGI detectors, because FPBA can deliver successful black-box attacks across models, generators, defense methods, and even evade cross-generator detection, which is a crucial real-world detection scenario. The code will be shared upon acceptance.
AI-generated Image Detection, Adversarial Examples
ccs: Security and privacy Software and application securityccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Computer vision
1. Introduction
The notable progress in generative models, such as GANs (Goodfellow et al., 2014a) and Diffusion models (Ho et al., 2020), is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
Consequently, numerous detectors have been proposed(Durall et al., 2020; Chai et al., 2020; Corvi et al., 2023b) to identify AI-generated images (AIGI). Recent state-of-the-art detectors (Wang et al., 2020; Zhong et al., 2023) rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Despite several works (Carlini and Farid, 2020; Hussain et al., 2021; Neekhara et al., 2021; Hou et al., 2023; Jia et al., 2022a) exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored. Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs. These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
Figure 1. Illustration of the proposed method.
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works (Dzanic et al., 2020; Frank et al., 2020) have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models. To further improve the adversarial transferability across heterogeneous models, e.g. transferring adversarial examples across Visual Transformers (ViTs) and Convolutional Neural Networks (CNNs), we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose the first adversarial attack for general AIGI detection, to add adversarial perturbations in the frequency domain from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
The contributions of this work can be summarized as follows:
  • To our best knowledge, we are the first to systemically investigate the adversarial vulnerability of state-of-the-art AIGI detectors. We extensively evaluate the adversarial robustness of traditionally trained models, defense models, and cross-generated detection within real-world scenarios. We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
  • We propose a novel attack against AIGI detection by exploring the vulnerable frequency region in a Bayesian manner. A new post-train Bayesian strategy for attack is proposed to explore the full posterior distribution over the surrogate, without the need for re-training the substitute model.
  • We conduct comprehensive experiments on eight spatial-based detectors with CNN-based, Vit-based and CLIP-based architectures, and two frequency-based detectors using AIGI and Deepfake datasets. The extensive results demonstrate our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
2. Related Work
AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs (Goodfellow et al., 2014a), VAEs (Kingma and Welling, 2014), and Diffusion models (Ho et al., 2020), has brought about substantial risks in the dissemination of disinformation. Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods to primarily focus on face forgeries. However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects. Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances. Data-driven-based approaches (Durall et al., 2020; Chai et al., 2020; Wang et al., 2020; Corvi et al., 2023b) have demonstrated promising results on generated data identically distributed. However, their performance significantly degraded when presented with previously unseen generated images. More recently, certain methods (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) focus on analyzing forgery traces in the frequency domain and learning relevant forgery features (Corvi et al., 2023a; Liu et al., 2023; Pontorno et al., 2024; Zhang et al., 2019) to enhance the generalization capability of the detection system. In pursuit of these objectives, some methods have been proposed to learn local forgery features (Ju et al., 2023; Lorenz et al., 2023) in the spatial domain. AIGI fingerprint representations (Jeong et al., 2022) are acquired through the learning of noise patterns. Notably, studies such as (Liu et al., 2022; Zhong et al., 2023) have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance. A sequence of detection methodologies (Tan et al., 2023; Liu et al., 2023; Cozzolino et al., 2023; Chang et al., 2023) has been progressively investigated, relying on the practice of freezing pre-trained models. These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers. For instance, UniFD (Ojha et al., 2023) explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection. Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error (Wang et al., 2023a; Ricker et al., 2024), adversarial teacher-student discrepancy-aware framework (Zhu et al., 2023a) or diffusion noise(Zhang and Xu, 2023).
The existing models consistently demonstrate enhancement in detecting AI-generated images. However, their behaviors under adversarial attacks have not been thoroughly examined. Our work is complementary to existing research by making the first attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly. The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including face recognition (Jia et al., 2022b), autonomous driving (Zhu et al., 2023b; Xie et al., 2023) and human activity recognition (Diao et al., 2021). Very recently, the adversarial attack against Deepfake detection has drawn attention. Hussain et al. (Hussain et al., 2021) employ gradient sign-based methods (Kurakin et al., 2018; Goodfellow et al., 2014b) to evaluate the adversarial robustness of video Deepfake detectors. Li et al. (Li et al., 2021) generate adversarial fake images along the latent face manifold. Carlini et al. (Carlini and Farid, 2020) and Neekhara et al. (Neekhara et al., 2021) investigate the adversarial threat to Deepfake detection under the black-box setting. Jia et al. (Jia et al., 2022a) propose to generate adversarial perturbations in the frequency domain. Hou et al. (Hou et al., 2023) narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations.
Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation. However, the adversarial robustness of AIGI detection has received little attention. In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
3. Methodology
Figure 2. The workflow of FPBA. We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real.
3.1. Preliminaries
Let $\mathbf{x}$ and $y$ represent the original image and its corresponding label. $f_{\Theta}$ denotes the AI-generated image detectors. We aim to inject adversarial perturbation into the original image that makes the detector misclassify. Such an adversary problem can be optimized by minimizing the predictive probability, i.e. maximizing the classification loss:
(1) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
where $L$ is the binary cross entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD (Madry et al., 2017) or I-FGSM (Kurakin et al., 2018):
(2) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
(3) $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$
3.2. Frequency-based Analysis and Attacks
[…]
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts (Wang et al., 2020; Jeong et al., 2022; Zhong et al., 2023). While these subtle clues are invisible in the spatial domain, a series of works (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map (Long et al., 2022) to visualize the sensitive components of real and fake images across different models:
(4) $\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})}$
[…]
(5) $\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$
where $\Gamma(\cdot)$ denotes the random spectrum transformation (Long et al., 2022). $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attack in the frequency domain via spectrum transformation:
(6) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
3.3. Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 6 can simulate different substitute models with a homogeneous architecture (Long et al., 2022), it shows limited transferability when applied to heterogeneous architectures, e.g. transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e. exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 6 by minimizing the Bayesian posterior predictive distribution:
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,p(y\mid\Gamma(\tilde{\mathbf{x}}),\mathcal{D})$
(7) $=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
where $p(\Theta\mid\mathcal{D})\propto p(\mathcal{D}\mid\Theta)p(\Theta)$. $\mathcal{D}$ is the dataset and $p(\Theta)$ is the prior of model weights. Attacking Bayesian Neural Networks (BNNs) rather than a single DNN allows for the output fusion from an ensemble of infinitely many DNNs with diverse predictions, thereby improving adversarial transferability.
Figure 4. The architecture of the appended model. $\sigma$ means the sigmoid layer.
3.3.1. Post-train Bayesian Strategy for Frequency-based Attack
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs (Izmailov et al., 2021). Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets (Zhu et al., 2024; He et al., 2021). From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
Therefore, we propose a post-train Bayesian strategy to turn a single surrogate into a Bayesian one, without the need for re-training. The parameters over the pre-trained surrogate are represented as $\Theta=[\theta,\theta_{c}]$, in which $f_{\theta}$ represents the feature extraction backbone, and $f_{\theta_{c}}$ represents the fully-connected layer network for classification. As shown in Fig. 4, we fix the pre-trained surrogate and append a tiny Bayesian component $g_{\theta^{\prime}}$ behind the feature extraction backbone $f_{\theta}$. The new logits can be computed via a skip connection:
(8) $\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x})$
We choose to apply Bayesian Model Averaging to optimize the appended Bayesian model:
(9) $\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)$
where $K$ is the number of appended models. $\Theta$ is fixed to avoid re-training. We surprisingly find that adopting a simple MLP layer for appended models works well in all cases, hence training the appended models is much faster than re-training a surrogate. Finally, the frequency-based post-train Bayesian attack can be conducted with iterative gradient-based methods:
(10) $\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}$
3.3.2. Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo(Springenberg et al., 2016) to sample appended model $\theta^{\prime}$ in each iteration:
$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I})$
(11) $\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2}$
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically (Springenberg et al., 2016).
3.4. Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain (Wang et al., 2020; Jeong et al., 2022). We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
(12) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\}$
(13) $g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i}$
(14) $d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k})$
where $g^{i}_{k}$ and $d^{i}_{k}$ are the gradients computed in the frequency domain and spatial domain respectively. For frequency gradient, we conduct random spectrum transformation with $N$ times to get more diverse spectrums. Our proposed method leverages both spatial attack gradients and frequency attack gradients in a Bayesian manner, aiming to further narrow the discrepancy between surrogate models and victim models. The complete algorithm of our method is presented in Algorithm 1. An overview illustration of FPBA is shown in Fig. 2.
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models; ;
Output: The adversarial example $\tilde{\mathbf{x}}$;
// Post-train Bayesian Optimization
Randomly initialize $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
for j = 1 to $N_{tra}$ do
       for n = 1 to $K$ do
             Randomly sample a mini-batch data $\{\mathbf{x},y\}_{j}$;
             Compute $\mathbf{h}_{\theta^{\prime}_{k}}=\frac{\partial logp(y|\mathbf{x},\Theta,\theta^{\prime}_{k})}{\partial\theta^{\prime}_{k}}$;
             for t = 1 to $M_{\theta^{\prime}}$ do
                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Sec. 3.3.2;
             end for
       end for
end for
return $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
// Frequency-based Post-train Bayesian Attack
$\tilde{\mathbf{x}}^{0}=\mathbf{x}$ ;
for i = 1 to $I$ do
       for n = 1 to $N$ do
             Get spectrum transformation output $\Gamma(\tilde{\mathbf{x}}^{i})$ using Eq. 5 ;
       end for
      for k = 1 to $K$ do
             Averaging frequency gradient $g_{k}$ using Eq. 13 ;
             Calculating spatial gradient $d_{k}$ using Eq. 14 ;
       end for
      Sample $\tilde{\mathbf{x}}^{i+1}$ from $\tilde{\mathbf{x}}^{i}$ via Eq. 12 ;
end for
return $\tilde{\mathbf{x}}$;
Algorithm 1 Inference on FPBA
4. Experimets
4.1. Experimental Settings
Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection. Synthetic LSUN is a commonly used dataset proposed by CNNSpot (Wang et al., 2020), containing 720k real images from LSUN and 360k fake images generated by ProGAN (Karras et al., 2018). GenImage (Zhu et al., 2024) is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in (Zhu et al., 2024), we employ a subset of GenImage collecting 162k real images from Imagenet (Deng et al., 2009) and 162k Stable Diffusion(SD) V1.4 (Rombach et al., 2022) generated images for training. The images generated by the other generators are used for testing in Sec. 4.3.2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by (Shamshad et al., 2023). The Deepfake dataset consists of 50k real face images from FFHQ (Karras et al., 2019) and 50k generated face images generated by StyleGAN2 (Karras et al., 2019). After training, we collect only the correctly classified testing samples for attack in evaluation.
Evaluated Models: We extensively evaluate the transferability of adversarial examples by eight different models across heterogeneous architectures under the spatial and frequency domain. We choose spatial-based classification models ResNet-50 (He et al., 2016) and MobileNet (Howard et al., 2017) as the source models. For target models, apart from the above models, we also use spatial-based models including EfficientNet (Tan and Le, 2019), Vision Transformer(Vit) (Dosovitskiy et al., 2020) and Swin-Transformer(Swin-Vit) (Liu et al., 2021), and frequency-based detectors DCTA (Frank et al., 2020) and Spec (Zhang et al., 2019). We report the accuracy of these detectors in the supplement. Besides, we also evaluate the robustness of state-of-the-art AIGI detection methods, including GramNet (Liu et al., 2020), LGrad (Tan et al., 2023), LNP (Liu et al., 2022) and UnivFD (Ojha et al., 2023).
Compared Methods: We adopt gradient-based methods, I-FGSM (Kurakin et al., 2018), PGD (Madry et al., 2017), MI (Dong et al., 2018) and Frequency-based methods S2(Long et al., 2022). We also consider ensemble-based methods Ensemble (Dong et al., 2018) and SVRE (Xiong et al., 2022). For a fair comparison, we ran 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all attacks.
Implementation Details: We choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001. We follow the same data augmentation strategy in CNNSpot (Wang et al., 2020). Before cropping, images are blurred with $\sigma$ Uniform[0, 3] with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels and apply ImageNet normalization. For post-train Bayesian optimization, we follow the default setting in (Wang et al., 2023b). Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following (Long et al., 2022). All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
Table 1. The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN, GenImage subset and Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
ResNet PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
MobileNet PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
IFGSM 78.3 71.5 30.9 34.0 27.1 28.3 11.0 27.8 32.9
MIFGSM 78.3 75.3 30.1 35.8 27.8 28.3 15.0 28.2 34.4
ResNet-50 PGD 99.3 88.0 48.5 52.2 48.6 49.9 15.4 48.1 50.1
S2I 98.0 91.0 66.8 64.7 36.7 50.1 16.9 53.0 54.2
FPBA(Ours) 100 96.6 59.4 67.3 47.9 49.8 21.1 51.2 56.2
IFGSM 9.0 8.9 8.9 58.9 9.1 9.1 4.4 8.5 8.3
MIFGSM 9.7 9.2 8.9 59.0 9.5 9.0 5.9 8.8 8.7
MobileNet PGD 42.3 39.9 36.4 97.1 48.4 49.8 14.2 40.0 38.7
S2I 35.6 33.1 26.8 78.7 24.1 43.4 6.2 28.0 28.2
GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
ResNet PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
MobileNet PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
4.2. The Evaluation on Normally Trained Models
We show the attack performance against all models on all datasets in Tab. 1. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.8%, while S2I, PGD, IFGSM and MIFGSM only have 95.6%, 95.4%, 65.4% and 65.5%. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 55.3%, surpassing MIFGSM, PGD and S2I by a margin of 27.7%, 7.6% and 8.4% respectively.
Next, We find that the attack results for FPBA are usually not the best on frequency-based models. To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq. 4 in Fig. 5. Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig. 3), the frequency components focused on in frequency-based models are very sparse. Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain. Nevertheless, our method still has the best average success rate over all models. In addition, we combine spatial-based gradient and frequency-based to further improve the attack performance on frequency-based models. FPBA gets an average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than SSA, MIFGSM and IFGSM respectively.
Finally, FPBA has nearly top-2 results in all 48 cases (datasets vs. victim models vs. target models vs. attack setting), which shows consistent attack performance across different attack settings (white-box vs. black-box), tasks (AIGI detection vs. Deepfake detection) and models (CNN-based vs. Vit-based vs. Frequency-based detectors). We leave the detailed robustness analysis across models to Sec. 4.4.1.
Figure 5. Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGan datasets) for frequency-based models. The frequency components of frequency-based models are highly sparse in comparison with spatial-based models.
Table 2. The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of ResNet-50, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of ResNet, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of ResNet, DCTA and ViT. 4 means only using ResNet as surrogate model.
Attack ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave
ENS1 55.9 45.4 69.9 76.2 16.8 42.9 23.3 43.7 34.4
SVRE1 63.7 64.5 75.5 72.9 10.3 30.4 28.6 64.2 39.6
ENS2 57.3 58.6 54.1 92.8 19.5 38.5 77.9 59.7 46.0
SVRE2 77.7 77.9 76.4 97.8 13.9 33.6 70.1 78.1 56.0
ENS3 9.7 6.3 4.8 10.1 37.0 72.9 20.6 5.5 12.7
SVRE3 22.6 12.9 8.1 13.2 29.4 87.0 24.0 9.6 14.6
FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
4.2.1. Comparsion with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods, including ENSEMBLE  (Dong et al., 2018) and SVRE (Xiong et al., 2022). Since there are three types of models for AIGI detection (CNN-based vs. Vit-based vs. Frequency-based detectors), we hence conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (ResNet-50, MobileNet, EfficientNet); (2) taking CNNs and Vits as ensemble surrogates (ResNet-50, MobileNet, Vit); (3) taking CNNs, Vits and frequency-based detectors as ensemble surrogates (ResNet-50, DCTA, Vit). Although we can also use more than one architecture for our method, we only use ResNet-50 as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Tab. 2. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across Vit-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on ResNet still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by ResNet-50 with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of ResNet-50, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
Table 3. Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot (Wang et al., 2020).
Attack GramNet LGrad LNP UnivFD Ave
IFGSM 39.4 7.5 37.0 14.6 24.6
MIFGSM 9.65 6.6 38.5 24.2 19.7
PGD 67.7 50.2 47.4 11.5 44.2
S2I 49.1 49.1 28.3 8.3 33.7
FPBA 87.8 50.2 39.5 16.6 48.5
4.3. The Evaluation on Diverse AIGI Detections
4.3.1. Attack SOTA AIGI Detectors
In this section, we examine the robustness of state-of-the-art (SOTA) AIGI detectors against adversarial attacks and present the results in Tab. 3. Even though these detectors can achieve SOTA detection performance on benign data, they still suffer from a serious adversarial threat and demonstrate our method is the strongest threat to these SOTA detectors. Unlike common detectors trained on the fake images for the task of binary classification, UnivFD (Ojha et al., 2023) extract fixed features via CLIP (Radford et al., 2021) and use nearest neighbor or linear probing for classification. The obvious diversity of the surrogate model(CNNSpot) and victim model(UnivFD) leads to a limited adversarial transferability on each other. Nevertheless, our method still achieves the best results on UnivFD compared to other attacks.
Table 4. Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
Detector Acc Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
All 60.5 100.0 99.9 50.8 99.7 50.0 76.8
Real 100.0 100.0 100.0 100.0 100.0 99.9 100.0
Swin-ViT Fake 21.0 100.0 99.9 1.5 99.3 0.1 53.6
All 62.0 99.6 99.5 50.3 97.9 49.8 76.5
Real 99.4 99.6 99.6 99.5 99.6 99.4 99.5
ResNet Fake 24.7 99.6 99.4 1.2 96.1 0.2 53.5
All 58.4 99.5 99.3 50.2 96.0 50.2 75.6
Real 99.5 99.6 99.5 99.5 99.5 99.3 99.5
ViT Fake 17.4 99.3 99.1 0.9 92.4 1.1 51.7
Table 5. The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-Vit.
Victim Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
Swin-ViT 97.6 97.4 97.7 96.2 96.9 97.1 97.1
ResNet 22.9 44.2 44.8 5.0 44.9 6.2 28.0
ViT 19.9 14.5 16.6 3.3 20.1 2.1 12.7
Table 6. The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-Vit.
Victim ASR Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
Real 94.9 96.4 96.2 96.0 96.0 97.0 95.9
Swin-ViT Fake 96.5 98.3 98.7 100.0 97.9 100.0 98.3
Real 3.0 3.4 2.9 2.0 3.6 4.6 3.2
ResNet Fake 100.0 85.1 86.9 100.0 86.3 100.0 93.0
Real 0.1 0.0 0.3 0.3 0.4 0.4 0.3
ViT Fake 96.1 29.0 32.9 100.0 39.8 100.0 66.3
4.3.2. Attack Cross-Generator Image Detection
One important real-world detection problem is cross-generator image detection, i.e. identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 (Rombach et al., 2022) and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney (Midjourney, 2022), SD V1.4 (Rombach et al., 2022), SD V1.5 (Rombach et al., 2022), ADM (Dhariwal and Nichol, 2021), Wukong (Wukong, 2022) and BigGAN (Brock et al., 2018). Because Swin-Vit achieves the SOTA results on different subsets (Zhu et al., 2024), we use it as the surrogate model.
The benign accuracy and attack performance on unseen source data are reported in Tab. 4 and Tab. 5 respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Tab. 4,Tab. 6), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
4.3.3. Attack Defense Models
Furthermore, we evaluate FPBA on attacking defense models. We first employ adversarial training (Madry et al., 2017) for evaluation, which is the most popular defense baseline. However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection. Considering that AIGI detection typically (Wang et al., 2020) uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from (Wang et al., 2020; Zhu et al., 2024) to utilize Gaussian blurring and JPEG compression as a defense. As shown in Tab. 7, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
Table 7. Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
Swin-ViT
Blur+JPEG(0.1) Blur+JPEG(0.5)
IFGSM 98.8 86.6(-12.2)
MIFGSM 98.3 92.7(-5.6)
PGD 99.3 87.3(-12)
SSA 100.0 80.9(-19.1)
FPBA 100.0 99.2(-0.5)
4.4. Additional Performance Analysis
4.4.1. The Phenomenon of Gradient Masking in AIGI Detectors
In Tab. 1, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate (Dong et al., 2018). It indicates the possibility of gradient masking (Athalye et al., 2018). To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by ResNet-50 on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 6 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
Most gradient components of adversarial examples generated by IFGSM (Fig. 6(a)) and MIFGSM (Fig. 6 (c)) tend to stabilize around zero, indicating the vanishing gradient leading to a limited transferability to IFGSM and MIFGSM. Because the only difference between IFGSM and PGD is that PGD randomly chooses the starting point within the $l_{\infty}$ constraint, we apply random initialization to MIFGSM and find the value of gradient component increase (Fig. 6 (d)), w.r.t. the average success rate increasing from 40.4% to 58.6% (still lower 10.6% than ours). This analysis formally demonstrates the phenomenon of gradient masking in AIGI detectors. Therefore, we advocate for future work for attacks on AIGI detectors to employ randomized-based strategies to circumvent the effect of gradient masking. Our proposed method conducts the spectrum transformation in the frequency domain and hence is also effective for gradient masking.
Figure 6. The gradient components of ResNet on adversarial examples generated by different attack methods. (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
4.4.2. Qualitative Analysis
Table 8. The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
[…]
PGD 30.00 33.51 0.88
FPBA(Ours) 16.08 36.26 0.94
Figure 7. Visual comparison with PGD. (a): the original image generated by diffusion models. (b): adversarial examples crafted by PGD. (c): adversarial examples crafted by FPBA. The image quality of the adversarial example crafted by our method is much closer to the original image.
In this study, we evaluated the image quality of the generated adversarial samples using PSNR and SSIM scores as metrics. We present the quantitative results of MSE, PSNR(db) and SSIM in Tab. 8. The image quality of our method is higher than PGD. As illustrated in Fig. 7, FPBA can generate more natural-looking adversarial examples in comparison with PGD.
4.4.3. Ablation Study
Table 9. Ablation Study on ProGAN dataset. The adversarial samples are crafted from ResNet50.
Attack DenNet MobNet Spec ViT
Spatial 71.5 34.0 27.1 11.0
Frequency 91.0 64.7 36.7 16.9
Spatial-frequency 96.2 66.2 47.9 21.0
We conduct an ablation study in Tab. 9 to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
5. Conclusion
In this paper, we fully investigate the robustness of AIGI detectors against adversarial attacks. For the task of AIGI detection, we propose a novel frequency-based post-train Bayesian attack, which explores the vulnerable regions in the frequency domain in a Bayesian manner. More broadly, our proposed post-train Bayesian attack strategy extends the Bayesian attack family. Extensive experiments have been conducted across models, generators and defense methods under both white-box and black-box settings. The following conclusions can be drawn from our investigation.
  • The current SOTA AIGI detectors are vulnerable to adversarial attacks, while the widely-used defense mechanism adversarial training cannot be directly applied to the AIGI detectors, which brings great challenges to build reliable and robust detectors.
  • AIGI detectors rely on gradient masking to mitigate the transferability of gradient-based attacks, but the effect of gradient masking can be easily circumvented by a simple random initialization strategy.
  • The attack success rate is positively associated with the detection accuracy. Therefore, evaluating detector robustness on low-accuracy subsets is futile, as there are no effective gradients for attack.
These findings indicate that building robust AIGI detectors is still an open problem. We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to the robustness of AIGI detectors.