### v3

# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

Abstract
Recent advancements in image synthesis, particularly with the advent of GAN and Diffusion models, have amplified public concerns regarding the dissemination of disinformation. To address such concerns, numerous AI-generated Image (AIGI) Detectors have been proposed and achieved promising performance in identifying fake images. However, there still lacks a systematic understanding of the adversarial robustness of AIGI detectors. In this paper, we examine the vulnerability of state-of-the-art AIGI detectors against adversarial attack under white-box and black-box settings, which has been rarely investigated so far. To this end, we propose a new method to attack AIGI detectors. First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution. Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous AIGI detectors, e.g. transferring adversarial examples across CNNs and ViTs. This is achieved by introducing a novel post-train Bayesian strategy that turns a single surrogate into a Bayesian one, capable of simulating diverse victim models using one pre-trained surrogate, without the need for re-training. We name our method as Frequency-based Post-train Bayesian Attack, or FPBA. Through FPBA, we show that adversarial attack is truly a real threat to AIGI detectors, because FPBA can deliver successful black-box attacks across models, generators, defense methods, and even evade cross-generator detection, which is a crucial real-world detection scenario. The code will be shared upon acceptance.
Index Terms:
AI-generated Image Detection, Adversarial Examples.
I Introduction
Figure 1: A high-level illustration of our proposed method.
The notable progress in generative models, such as GANs [1] and Diffusion models [2], is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI). Recent state-of-the-art detectors [3, 4] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Despite several works [5, 6, 7, 8, 9] exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored. Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs. These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works [10, 11] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains. Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[12], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g. transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table I). To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
The contributions of this work can be summarized as follows:
  • We systemically investigate the adversarial robustness of state-of-the-art AIGI detectors, including standard trained models, defense models, and cross-generated detection within real-world scenarios. We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
  • We propose a novel attack against AIGI detection by exploring the vulnerable frequency region in a Bayesian manner. A new post-train Bayesian strategy for attack is proposed to explore the full posterior distribution over the surrogate, without the need for re-training the substitute model.
  • We conduct comprehensive experiments on 13 various AIGI detectors, including heterogeneous model architectures and various detection methods. Extensive results demonstrate that our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
II Related Work
AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs [1], VAEs [13], and Diffusion models [2], has brought about substantial risks in the dissemination of disinformation. Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods [14, 15, 16], to primarily focus on face forgeries. However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects. Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances. Data-driven-based approaches [3, 17] have demonstrated promising results on generated data identically distributed. However, their performance significantly degraded when presented with previously unseen generated images. More recently, certain methods [10, 11] focus on analyzing forgery traces in the frequency domain and learning relevant forgery features [18] to enhance the generalization capability of the detection system. In pursuit of these objectives, some methods have been proposed to learn local forgery features [19, 20] in the spatial domain. AIGI fingerprint representations [21] are acquired through the learning of noise patterns. Notably, studies such as [22, 4] have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance. A sequence of detection methodologies [23] has been progressively investigated, relying on the practice of freezing pre-trained models. These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers. For instance, UniFD [24] explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection. Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error [25, 26], adversarial teacher-student discrepancy-aware framework [27] or diffusion noise[28]. The existing models consistently demonstrate enhancement in detecting AI-generated images. However, their behaviors under adversarial attacks have not been thoroughly examined. Our work is complementary to existing research by making the attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly. The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including image classification[29, 30], multimedia communication [31] and human activity recognition [32]. Very recently, the adversarial attack against Deepfake detection has drawn attention. Hussain et al. [6] employ gradient sign-based methods [33, 34] to evaluate the adversarial robustness of video Deepfake detectors. Li et al. [35] generate adversarial fake images along the latent face manifold. Carlini et al. [5] and Neekhara et al. [7] investigate the adversarial threat to Deepfake detection under the black-box setting. Jia et al. [9] propose to generate adversarial perturbations in the frequency domain. Hou et al. [8] narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations. FakePolisher [36] and TraceEvader [37] improve the untraceability of deepfakes by removing artificial fingerprints. Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation. However, the adversarial robustness of AIGI detection has received little attention. In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
III Methodology
Figure 2: The workflow of FPBA. We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real. DCT and IDCT are discrete cosine transformation and inverse discrete cosine transformation respectively.
III-A Preliminaries
Let $\mathbf{x}$ and $y$ represent the original image and its corresponding label. $f_{\Theta}$ denotes the AI-generated image detectors. We aim to inject adversarial perturbation into the original image that makes the detector misclassify. Such an adversary problem can be optimized by minimizing the predictive probability, i.e. maximizing the classification loss:
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon,$ (1)
where $L$ is the binary cross entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD [38] or I-FGSM [33]:
$\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
$=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$ (2)
[…]
III-B Frequency-based Analysis and Attacks
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 21, 4]. While these subtle clues are invisible in the spatial domain, a series of works [10, 11] demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [39] to visualize the sensitive components of real and fake images across different models:
$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
[…]
$\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
where $\Gamma(\cdot)$ denotes the random spectrum transformation [39]. $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attack in the frequency domain via spectrum transformation:
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon.$ (5)
III-C Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 5 can simulate different substitute models with a homogeneous architecture [39], it shows limited transferability when applied to heterogeneous architectures, e.g. transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e. exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 5 by minimizing the Bayesian posterior predictive distribution:
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,p(y\mid\Gamma(\tilde{\mathbf{x}}),\mathcal{D})$
$=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon,$ (6)
where $p(\Theta\mid\mathcal{D})\propto p(\mathcal{D}\mid\Theta)p(\Theta)$. $\mathcal{D}$ is the dataset and $p(\Theta)$ is the prior of model weights. Attacking Bayesian Neural Networks (BNNs) rather than a single DNN allows for the output fusion from an ensemble of infinitely many DNNs with diverse predictions, thereby improving adversarial transferability.
Figure 4: The architecture of the appended model. $\sigma$ means the sigmoid layer.
III-C1 Post-train Bayesian Strategy
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs [40]. Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [12, 41]. From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
Therefore, we propose a post-train Bayesian strategy to turn a single surrogate into a Bayesian one, without the need for re-training. The parameters over the pre-trained surrogate are represented as $\Theta=[\theta,\theta_{c}]$, in which $f_{\theta}$ represents the feature extraction backbone, and $f_{\theta_{c}}$ represents the fully-connected layer network for classification. As shown in Fig. 4, we fix the pre-trained surrogate and append a tiny Bayesian component $g_{\theta^{\prime}}$ behind the feature extraction backbone $f_{\theta}$. The new logits can be computed via a skip connection:
$\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x}).$ (7)
We choose to apply Bayesian Model Averaging to optimize the appended Bayesian model:
$\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),$
$\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta),$ (8)
where $K$ is the number of appended models. $\Theta$ is fixed to avoid re-training. We surprisingly find that adopting a simple MLP layer for appended models works well in all cases, hence training the appended models is much faster than re-training a surrogate. Finally, the frequency-based post-train Bayesian attack can be conducted with iterative gradient-based methods:
$\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}.$ (9)
III-C2 Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[42] to sample appended model $\theta^{\prime}$ in each iteration:
$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I}),$
$\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2},$ (10)
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [42].
III-D Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3, 21]. We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
$\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\},$ (11)
$g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i},$ (12)
$d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k}),$ (13)
where $g^{i}_{k}$ and $d^{i}_{k}$ are the gradients computed in the frequency domain and spatial domain respectively. For frequency gradient, we conduct random spectrum transformation with $N$ times to get more diverse spectrums. Our proposed method leverages both spatial attack gradients and frequency attack gradients in a Bayesian manner, aiming to further narrow the discrepancy between surrogate models and victim models. The complete algorithm of our method is presented in Algorithm 1. An overview illustration of FPBA is shown in Fig. 2.
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
Output: The adversarial example $\tilde{\mathbf{x}}$;
// Post-train Bayesian Optimization
Randomly initialize $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
for j = 1 to $N_{tra}$ do
       for n = 1 to $K$ do
             Randomly sample a mini-batch data $\{\mathbf{x},y\}_{j}$;
             Compute $\mathbf{h}_{\theta^{\prime}_{k}}=\frac{\partial logp(y|\mathbf{x},\Theta,\theta^{\prime}_{k})}{\partial\theta^{\prime}_{k}}$;
             for t = 1 to $M_{\theta^{\prime}}$ do
                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Section III-C2;
             end for
       end for
end for
return $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
// Frequency-based Post-train Bayesian Attack
$\tilde{\mathbf{x}}^{0}=\mathbf{x}$ ;
for i = 1 to $I$ do
       for n = 1 to $N$ do
             Get spectrum transformation output $\Gamma(\tilde{\mathbf{x}}^{i})$ using Eq. 4 ;
       end for
      for k = 1 to $K$ do
             Average frequency gradient $g_{k}$ using Eq. 12 ;
             Calculate spatial gradient $d_{k}$ using Eq. 13 ;
       end for
      Sample $\tilde{\mathbf{x}}^{i+1}$ from $\tilde{\mathbf{x}}^{i}$ via Eq. 11 ;
end for
return $\tilde{\mathbf{x}}$;
Algorithm 1 Inference on FPBA
IV Experimets
IV-A Experimental Settings
Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection. Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43]. GenImage [12] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in [12], we employ a subset of GenImage collecting 162k real images from Imagenet [44] and 162k Stable Diffusion(SD) V1.4 [45] generated images for training. The images generated by the other generators are used for testing in Section IV-C2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [46]. The Deepfake dataset consists of 50k real face images from FFHQ [47] and 50k generated face images generated by StyleGAN2 [47]. After training, we collect only the correctly classified testing samples for attack in evaluation.
Evaluated Models: We extensively evaluate the transferability of adversarial examples by 13 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods. For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [48], EfficientNet [49] and DenseNet [50], and ViT-based detectors Vision Transformer(ViT) [51] and Swin-Transformer(Swin-ViT) [52]. For evaluating on various detection methods, we use frequency-based detectors DCTA [11] and Spec [18], gradient-based detectors LGrad [23], CLIP-based detectors UnivFD [24], and diffusion-based detectors DNF [28]. In addition, LNP [22] extracts the noise pattern of images and GramNet [53] learns the global texture representation. We also consider them as victim models.
Compared Methods: We adopt gradient-based methods, I-FGSM [33], PGD [38], MI [54] and frequency-based methods S2[39], SSAH [55] and ensemble-based methods ENS [54] and SVRE [56]. Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks Fakepolisher [36] and TraceEvader [37] as baselines. Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings. For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3]. Specificaly, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001. For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness. Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels on LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3]. On the FFHQ (StyleGAN2) dataset we resized images to 224 pixels to ensure the integrity of the real/fake faces. Subsequently, we apply ImageNet normalization across three datasets. For other AIGI detectors, we use the pre-trained model from their official code. For post-train Bayesian optimization, we follow the default setting in [57]. Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [39]. All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
TABLE I: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
Surrogate Model Attack Methods CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Average
IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
PGD 78.3 77.6 49.4 73.0 25.6 47.6 41.0 70.2 54.9
CNNSpot S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
SSAH 97.8 2.5 1.3 3.0 0.8 1.0 1.0 0.7 1.4
FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
MIFGSM 23.8 23.9 28.6 75.4 19.6 42.0 15.5 24.8 25.5
PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
MobileNet S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
SSAH 0.6 1.0 1.5 99.2 0.5 0.8 1.0 1.1 0.9
LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
IFGSM 78.3 71.5 30.9 34.0 27.1 28.3 11.0 27.8 32.9
MIFGSM 78.3 75.3 30.1 35.8 27.8 28.3 15.0 28.2 34.4
PGD 99.3 88.0 48.5 52.2 48.6 49.9 15.4 48.1 50.1
CNNSpot S2I 98.0 91.0 66.8 64.7 36.7 50.1 16.9 53.0 54.2
SSAH 94.8 5.4 0.5 1.1 0.6 3.6 0.7 0.3 1.7
FPBA(Ours) 100 96.6 59.4 67.3 47.9 49.8 21.1 51.2 56.2
IFGSM 9.0 8.9 8.9 58.9 9.1 9.1 4.4 8.5 8.3
MIFGSM 9.7 9.2 8.9 59.0 9.5 9.0 5.9 8.8 8.7
PGD 42.3 39.9 36.4 97.1 48.4 49.8 14.2 40.0 38.7
MobileNet S2I 35.6 33.1 26.8 78.7 24.1 43.4 6.2 28.0 28.2
SSAH 1.1 2.0 0.6 96.7 0.5 1.9 1.2 2.2 1.4
GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
IV-B Attack on Spatial-based and Frequency-based Detectors
We report the attack performance against spatial-based and frequency-based detectors in Table I. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.6%, while S2I, PGD, IFGSM and MIFGSM only have 93.1%, 93.1%, 66.2% and 66.2%. Although SSAH has a relatively high white-box success rate (97.7%), it fails to transfer the adversarial examples to black-box models. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 50.5%, surpassing the IFGSM, MIFGSM, PGD, S2I and SSAH by a big margin of 25.7%, 23.3%, 8.4%, 11.9% and 49.1%. Note that our proposed FPBA significantly outperforms the SOTA frequency-based attack S2I and SSAH, demonstrating that FPBA is a stronger frequency-based attack against AIGI detectors.
TABLE II: The attack success rate(%) compared with attacks agaisnt Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.
Datasets Attack CNNSpot DenNet EffNet MobNet ViT Swin Ave.
FakePolisher 94.5 98.3 71.6 88.6 15.1 99.5 77.9
LSUN TraceEvader 0.6 1.3 4.5 4.8 8.10 11.4 1.9
(ProGAN) FPBA 100.0 100.0 98.2 99.5 38.6 99.9 89.4
FakePolisher 93.5 93.2 12.2 95.5 33.5 97.6 70.9
GenImage TraceEvader 46.1 34.3 9.4 37.0 15.2 44.3 31.1
(SD) FPBA 100.0 99.7 97.7 98.9 42.1 97.9 89.4
Figure 5: Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGAN datasets) for frequency-based models. The frequency components of frequency-based models are highly sparse in comparison with spatial-based models. The color value represents the absolute gradient value of the model loss function after max-min normalization.
IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE III: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of CNNSpot, DCTA and ViT. 4 means only using CNNSpot as surrogate model.
Attack CNNSpot DenNet EffNet MobNet Spec DCTA ViT Swin Ave
ENS1 55.9 45.4 69.9 76.2 16.8 42.9 23.3 43.7 34.4
SVRE1 63.7 64.5 75.5 72.9 10.3 30.4 28.6 64.2 39.6
ENS2 57.3 58.6 54.1 92.8 19.5 38.5 77.9 59.7 46.0
SVRE2 77.7 77.9 76.4 97.8 13.9 33.6 70.1 78.1 56.0
ENS3 9.7 6.3 4.8 10.1 37.0 72.9 20.6 5.5 12.7
SVRE3 22.6 12.9 8.1 13.2 29.4 87.0 24.0 9.6 14.6
FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
TABLE IV: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
Attack GramNet LGrad LNP UnivFD DNF Ave
MIFGSM 9.7 6.6 38.5 24.2 46.7 25.1
PGD 67.7 50.2 47.4 11.5 48.3 45.0
S2I 49.1 49.1 28.3 8.3 43.0 35.6
FPBA 87.8 50.2 39.5 16.6 49.8 48.9
TABLE V: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
Detector Acc Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
All 60.5 100.0 99.9 50.8 99.7 50.0 76.8
Real 100.0 100.0 100.0 100.0 100.0 99.9 100.0
Swin-ViT Fake 21.0 100.0 99.9 1.5 99.3 0.1 53.6
All 62.0 99.6 99.5 50.3 97.9 49.8 76.5
Real 99.4 99.6 99.6 99.5 99.6 99.4 99.5
CNNSpot Fake 24.7 99.6 99.4 1.2 96.1 0.2 53.5
All 58.4 99.5 99.3 50.2 96.0 50.2 75.6
Real 99.5 99.6 99.5 99.5 99.5 99.3 99.5
ViT Fake 17.4 99.3 99.1 0.9 92.4 1.1 51.7
Attacks against Deepfake detectors are similar to our attacks. However, Deepfake mainly manipulating or synthesizing faces, often crafted by GANs. In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models. The content includes humans, non-human entities, abstract art, landscapes, and more. These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors. For example, techniques like searching for adversarial points on the face manifold [35] or removing facial forgery traces [36, 37] may not be applicable for AIGI detectors. To demonstrate this, we compare with FakePolisher [36] and TraceEvader [37], which are both the SOTA attacks for Deepfake detection. Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images. As shown in Table II, FPBA significantly outperforms FakePolisher and TraceEvader on both AIGI datasets. Although FakePolisher performs better than TraceEvader, but the former largely reduces the image quality, which is very visible and raises suspicion (Fig. 7). We suspect this is because FakePolisher has only learned the face manifold, limiting its ability to reconstruct fine-grained details in AIGI images.
IV-B2 Comparison with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [54] and SVRE [56], i.e. utilizing an ensemble of surrogate models to generate adversarial examples. To examine the impact of different surrogate combinations(CNN-based vs. Vit-based vs. Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT). Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Table III. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by CNNSpot with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
IV-B3 Evaluation on Frequency-based Detectors
Although FPBA has the best average success rate, we find that the attack results are usually not the best on frequency-based models. To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq. 3 in Fig. 5. Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig. 3), the frequency components focused on in frequency-based models are very sparse. Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain. Nevertheless, FPBA still gets an top-2 average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than S2I, MIFGSM and IFGSM respectively.
TABLE VI: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
Victim Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
Swin-ViT 97.6 97.4 97.7 96.2 96.9 97.1 97.1
CNNSpot 22.9 44.2 44.8 5.0 44.9 6.2 28.0
ViT 19.9 14.5 16.6 3.3 20.1 2.1 12.7
TABLE VII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
Victim ASR Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
Real 94.9 96.4 96.2 96.0 96.0 97.0 95.9
Swin-ViT Fake 96.5 98.3 98.7 100.0 97.9 100.0 98.3
Real 3.0 3.4 2.9 2.0 3.6 4.6 3.2
CNNSpot Fake 100.0 85.1 86.9 100.0 86.3 100.0 93.0
Real 0.1 0.0 0.3 0.3 0.4 0.4 0.3
ViT Fake 96.1 29.0 32.9 100.0 39.8 100.0 66.3
IV-C Evaluation on Diverse Detection Strategies
IV-C1 Attack Various SOTA Detection Methods
Except for detecting in spatial and frequency domian using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts. Specifically, GramNet [53] extracts the global texture representation; UnivFD [24] captures semantic feature using pre-trained CLIP [58]; LNP [22] extracts the noise pattern of images; LGrad [23] extract gradient information through a pre-trained model and DNF [28] leverages diffusion noise features, extracting them through an inverse diffusion process [2]. These significant differences among various detection methods present challenges for the transferability of adversarial examples. To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table IV. Overall, the transfer success rate is relatively low compared to the results in Table I, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other. Nevertheless, FPBA still give the best adversarial transferability.
IV-C2 Attack Cross-Generator Image Detection
One important real-world detection problem is cross-generator image detection, i.e. identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 [45] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [59], SD V1.4 [45], SD V1.5 [45], ADM [60], Wukong [61] and BigGAN [62]. Because Swin-ViT achieves the SOTA results on different subsets [12], we use it as the surrogate model.
The benign accuracy and attack performance on unseen source data are reported in Table V and Table VI respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Table V,Table VII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
IV-D Attack Performance against Compression in Real Scenarios
Figure 6: Attack performace on ResNet, Efficient, MobileNet and Swin against JPEG compression on LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
In real-world scenarios, images are often compressed during propagation, which may also distort adversarial noise patterns. Therefore, ensuring the effectiveness of adversarial attacks on compressed images is crucial for simulating real-world scenarios. To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images. We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer. All adversarial samples are crafted from ResNet-50. The results are reported in Fig. 6. For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance. For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases. However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
TABLE VIII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
Swin-ViT
Blur+JPEG(0.1) Blur+JPEG(0.5)
IFGSM 98.8 86.6(-12.2)
MIFGSM 98.3 92.7(-5.6)
PGD 99.3 87.3(-12)
S2I 100.0 80.9(-19.1)
FPBA 100.0 99.2(-0.8)
IV-D1 Attack Defense Models
Furthermore, we evaluate FPBA on attacking defense models. We first employ adversarial training [38] for evaluation, which is the most popular defense baseline. However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection. Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 12] to utilize Gaussian blurring and JPEG compression as a defense. As shown in Table VIII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
TABLE IX: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
Surrogate Model Attack Methods CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Average
IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
CNNSpot S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
SSAH 99.0 54.5 44.1 41.0 2.6 7.2 1.0 4.2 22.1
FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
MobileNet S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
SSAH 1.9 36.4 43.7 98.5 2.0 6.2 0.8 2.4 13.3
FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
IV-E Visual Quality Analysis
TABLE X: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
[…]
PGD 30.00 33.51 0.88
FakePolisher 34.63 32.92 0.88
FPBA(Ours) 16.08 36.26 0.94
Figure 7: Visual comparison with PGD. (a): the original image generated by diffusion models. (b): adversarial examples crafted by PGD. (c): the original image generated by FakePolisher. (d): adversarial examples crafted by FPBA. The image quality of the adversarial example crafted by our method is much closer to the original image.
To demonstrate the superior image quality achieved by our method, we conduct both qualitative and quantitative assessments. We first visualize adversarial examples generated by PGD, FakePolisher and our proposed FPBA in Fig. 7. Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in. In addition, the fine-gradined details of object generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images. In contrast, FPBA generate more natural-looking adversarial examples. Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM. As reported in Table X, FPBA outperforms other baselines across all quality assessment by a large margin. This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
IV-F Additional Performance Analysis
IV-F1 The Phenomenon of Gradient Masking in AIGI Detectors
In Table I, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [54]. It indicates the possibility of gradient masking [63]. To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by CNNSpot on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 8 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
Most gradient components of adversarial examples generated by IFGSM (Fig. 8(a)) and MIFGSM (Fig. 8 (c)) tend to stabilize around zero, indicating the vanishing gradient leading to a limited transferability to IFGSM and MIFGSM. Because the only difference between IFGSM and PGD is that PGD randomly chooses the starting point within the $l_{\infty}$ constraint, we apply random initialization to MIFGSM and find the value of gradient component increase (Fig. 8 (d)), w.r.t. the average success rate increasing from 40.4% to 58.6% (still lower 10.6% than ours). This analysis formally demonstrates the phenomenon of gradient masking in AIGI detectors. Therefore, we advocate for future work for attacks on AIGI detectors to employ randomized-based strategies to circumvent the effect of gradient masking. Our proposed method conducts the spectrum transformation in the frequency domain and hence is also effective for gradient masking.
Figure 8: The gradient components of CNNSpot on adversarial examples generated by different attack methods. (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
IV-F2 Evaluation on Deepfake Datasets
As we mention before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [47]. We report the results in Table IX. Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
TABLE XI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
Attack DenseNet MobileNet Spec ViT
Spatial 71.5 34.0 27.1 11.0
Frequency 91.0 64.7 36.7 16.9
Spatial-frequency 96.2 66.2 47.9 21.0
TABLE XII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
$N$ ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
0 64.5 64.0 37.3 60.6 17.5 48.4 40.5 59.7 48.8
5 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
10 99.6 99.1 78.6 97.5 19.9 48.0 52.4 96.5 70.3
15 99.4 99.0 78.2 96.6 20.2 48.3 53.0 96.2 70.2
20 99.2 98.7 77.5 96.7 20.7 48.6 51.1 95.6 69.8
IV-G Ablation Study
Spatial and Frequency Attack
We conduct an ablation study in Table XI to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
Number ($N$) of Spectrum Transformations
In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models. The adversarial examples were crafted from ResNet-50 in LSUN(ProGAN) dataset. With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper. As shown in Table XII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate. Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation. Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.
V Conclusion
In this paper, we fully investigate the robustness of AIGI detectors against adversarial attacks. For the task of AIGI detection, we propose a novel frequency-based post-train Bayesian attack, which explores the vulnerable regions in the frequency domain in a Bayesian manner. More broadly, our proposed post-train Bayesian attack strategy extends the Bayesian attack family. Extensive experiments have been conducted across models, generators and defense methods under both white-box and black-box settings. The following conclusions can be drawn from our investigation.
  • The current SOTA AIGI detectors are vulnerable to adversarial attacks, while the widely-used defense mechanism adversarial training cannot be directly applied to the AIGI detectors, which brings great challenges to build reliable and robust detectors.
  • AIGI detectors rely on gradient masking to mitigate the transferability of gradient-based attacks, but the effect of gradient masking can be easily circumvented by a simple random initialization strategy.
  • The attack success rate is positively associated with the detection accuracy. Therefore, evaluating detector robustness on low-accuracy subsets is futile, as there are no effective gradients for attack.
These findings indicate that building an adversarial robust AIGI detector is still an open problem. We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to this domain.
Acknowledgments
This work was supported in part by the NSF China (No. 62302139) and FRFCU-HFUT (JZ2023HGTA0202, JZ2023HGQA0101).